Subject: Math-using addition to solve a system of linear equations
1. x+y=12
x-y=4

Answer by jim_thompson5910(35256)

Hint: Add the two equations to get


    x + y = 12
+   x - y = 4
---------------
   2x + 0y = 16



So we now have the equation 2x = 16. Solve this for x. Once you have the solution for x, use it to find y.